Seven more than 3 times a number is -5. find the number.
A number divided by 3, then increased by -2 is 10. What is the number.

Found 2 solutions by Here2Help, lmk1213:
Answer by Here2Help(19)

3x+7=-5
3x+7-7=-5-7
3x=-12
3x/3=-12/3
x=-4
Second problem:
n/3+(-2)=10
n/3-2+2=10+2
n/3=12/1
1n/1=36/1
n=36

Answer by lmk1213(1)

3x+7=-5
-7 -7
____________
3x/3=-12/3
x=-4